Find the pH of solution formed by dissolving 0.1 mol of acetic acid (Ka = 1.8 * 10-8) and 0.2 mol of sodium acetate in total volume of 1L.
1
Expert's answer
2013-01-28T09:44:00-0500
Ka of acetic acid is 1.8 x 10^-5 (not 1.8 x 10^-8). It's an acidic buffer

pH = pKa + log (S/A)

where S = concentration of the salt
A = concentration of the acid

Ka = 1.8 x 10^-5 pKa = -log Ka = -log (1.8 x 10^-5) = 4.74

S = no. of moles of salt / volume of the solution = 0.2 / 1 = 0.2 M
A = no. of moles of acid / volume of the solution = 0.1 / 1 = 0.1 M

So, pH = 4.74 + log (0.2/0.1) = 4.74 + log 2 = 4.74 + 0.30 = 5.04